OpenStudy (automaticloveletter):

The incorrect work of a student to solve an equation 2(y + 8) = 4y is shown below: Step 1: 2(y + 8) = 4y Step 2: 2y + 10 = 4y Step 3: 2y = 10 Step 4: y = 5 Which of the following explains how to correct Step 2 and shows the correct value of y? The equation should be y + 8 = 4y after division by 2; y = 4 The equation should be y + 8 = 4y after division by 2; y = 6 2 should be distributed as 2y + 16; y = 4 2 should be distributed as 2y + 16; y = 8

OpenStudy (automaticloveletter):

@TheSmartOne @Kim21

OpenStudy (anonymous):

Either A or B, since the person screwed up when they added 2+6 rather than multiplying 2*6 :b

OpenStudy (anonymous):

After that all you need to do is solve for y :)

OpenStudy (anonymous):

Step 2 SHOULD have been: 2y+12=4y

OpenStudy (anonymous):

2y+12=4y then subtract 2y on both sides 12=2y and then divide by 2 on both sides 6=y

OpenStudy (anonymous):

A. 2 should be distributed as 2y + 12; y = 6 ^this would be the correct answer...again so sorry that I was offline! If you were to double check, it 6 would work. If you ever get stuck on a problem like this, just plug in the y to see if the solution would be equal: - 2(y + 6) = 4y simplifies to 12+12=24

OpenStudy (anonymous):

thats all you have to do then since its then Step 2 :) @automaticloveletter